A firm can sell at a price of $100 per unit all of a particular commodity it produces. If x units are produced each day, the number of dollars in the total cost of each day's production is x^2+20x+700. How many units should be produced each day so that the firm is guaranteed a profit?
Answer by mananth(16946)

Cost of production of x units = x^2+20x+700.
sales from 100 units = 100x
..
BREAK EVEN
when both are same
x^2+20x+700= 100x
x^2-80x+700=0
x^2-70x-10x+700=0
x(x-70)-10(x-70)=0
(x-70)(x-10)=0
x= 10 OR 70
..
More than 10 units & less than 70 units must be manufactured to make a profit
....
m.ananth@hotmail.ca